OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 1 less than or greater to 12 or 2x greater than or equal to 30. A number line with an open circle on 13, shading to the left, and an open circle on 15, shading to the right A number line with a closed circle on 13, shading to the left, and a closed circle on 15, shading to the right A number line with an open circle on 13, an open circle on 15, and shading in between A number line with a closed circle on 13, a closed circle on 15, and shading in between

OpenStudy (anonymous):

@Aureyliant

OpenStudy (anonymous):

\[x-1\le 12\] \[x \le12+1~or~x \le 13\]  \[(-\infty,13]\cup[15,\infty )\]

OpenStudy (anonymous):

^

OpenStudy (anonymous):



OpenStudy (anonymous):

So it would be, A number line with a closed circle on 13, shading to the left, and a closed circle on 15, shading to the right ?

OpenStudy (anonymous):

no,it is B

OpenStudy (anonymous):

Lol, that's what I chose.